A(g)⇌B(g)+C(g) Find the equilibrium concentrations of A, B, and C for each of the following different values of Kc. Assume that the initial concentration of A in each case is 1.0 M and that the reaction mixture initially contains no products. Make any appropriate simplifying assumptions.

Kc= 2.0

Kc=Kc= 1.6×10−2

Kc=Kc= 1.6×10−5
Answer #1

А) - ВЯ? * 1.OM + НО ПО + tu + – № E (1.0-W) M NM x M к. - - CB][C] [А] - (1.0 - 3) б) к = 20 х = 2.0 1.o- * + 2.0 - 2.0 = 0il Kc = 1.68102 1.6x102 = x² 1.0-a x² + (1.6x10²) - 1.68100 =0 2 -0.1187 in At equillisium 0.8913 M. [A] = 1.0 -0.1187 = [B]